OpenStudy (anonymous):

If I know some of the lengths of a right triangle can I find the two missing angle measures?

OpenStudy (amistre64):

trig functions help sine, cosine, and tangent

OpenStudy (amistre64):

 \[\sin(\alpha)=\frac{y}{r}\] \[\cos(\alpha)=\frac{x}{r}\] \[\tan(\alpha)=\frac{y}{x}\]

OpenStudy (anonymous):

thank you I needed to refresh my memory!

OpenStudy (amistre64):

youre welcome